OpenStudy (anonymous):

Help with my math please?(: *Write an equation for each problem. Then solve the equation* 1. Fifteen added to the product of a number and 6. The result is 9. Find the number. 2. During an evening out, Dean paid a cab driver 20$. He then spent 25$ on dinner and half of his remaining money on a painting. H bought an umbrella for $23.75 and had $42.15 left. How much money did Dead have at the beginning of the week?

OpenStudy (anonymous):

6n +15=9

OpenStudy (anonymous):

solve for n: 6n+15 - 15 = 9-15 6n = -6 n=-1